I'm given a utility function U=xy and MUx=y while MUy=x. Price of x is $2 and the income is $40. To maximize utility subject to the budget constraint, 5 units of y is purchased.

QUESTION: What must be the price of y and the amount of x consumed?


Thank you :)


1
Expert's answer
2021-05-05T07:28:22-0400

"U=xy" , "MUx=y" ,"MUy=x"  "I=40"


At utility maximisation: "\frac{MUx}{MUy}=\frac{Px}{Py}"

"\frac{y}{x}=\frac{2}{Py}"

"2x=Py\times y"

"x=\frac
{Py\times 
y}{2}"


Budget constraint: "Px\times x+Py\times y=I"

"2x+Py\times y=40" and "y=5"

"\frac{2(5\times Py)}{2}+Py\times 5=40"


"10Py=\$40"


"Py=\$4"


Calculating for the amount of x consumed:

"x=\frac{\$4\times 5}{\$2}"


"x=10 units"